Four years ago, Bling Diamond, Inc., paid a dividend of $1.85 per share. The company paid a dividend of $2.27 per share yesterday. Dividends will grow over the next five years at the same rate they grew over the last four years. Thereafter, dividends will grow at 7 percent per year.

  

What will the company’s cash dividend be in seven years?
Answer #1

Company's cash dividend in 7 years will be $ 3.36

Step-1:Calculation of dividend growth rate in last 4 years
A = P*(1+g)^n Where,
2.27 = 1.85*(1+g)^4 A Current dividend
1.227027 = (1+g)^4 P Dividend of 4 years ago
1.227027 ^(1/4) = 1+g g Dividend growth rate
1.052479 = 1+g n Years of time
0.052479 = g
So, Growth rate of dividend over last 4 years is 5.25%
Step-2:Calculation of dividend in 7 years from now
D7 = D0*(1+g1)^5*(1+g2)^2 Where,
= 2.27*(1+0.0525)^5*(1+0.07)^2 D7 = Dividend 7 years from today = ?
= $       3.36 D0 = Current dividend = $       2.27
g1 = Growth rate over next 5 years = 5.25%
g2 = Growth rate 2 years therafter = 7%